OpenStudy (anonymous):

radical expressions please help me step by step on how to do this problems please I have to simplify and assume that the variables are real numbers ^3pie27(x+3)^3

OpenStudy (anonymous):

\[\sqrt[3]{27 (x+3)^3}\]

OpenStudy (anonymous):

does anyone know the radical expression

OpenStudy (asnaseer):

ok, firstly, do you know what this equals?\[\sqrt[3]{x^3}=?\]

OpenStudy (anonymous):

thnaks so much I have been reading this over and over and I am so confused on this week work I have 3 questions to go to be done with this homework and it has taken me forever

OpenStudy (anonymous):

x^6 right

OpenStudy (asnaseer):

no that would be \((x^3)^2\)

OpenStudy (asnaseer):

here we are finding the cube root of x cubed

OpenStudy (anonymous):

ok now you got me lost

OpenStudy (asnaseer):

ok - lets start with something simpler, what would this equal?\[\sqrt{x^2}=?\]

OpenStudy (anonymous):

so would that be (x^2)

OpenStudy (anonymous):

or x(x+1)

OpenStudy (asnaseer):

ok - it looks like you first need to learn about the various rules for exponents. I suggest you take a look here: http://www.mathsisfun.com/algebra/exponent-fractional.html and then come back to this problem.

OpenStudy (anonymous):

k I will look thanks

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

so that would mean that it would be (xx)

OpenStudy (asnaseer):

what would this equal?\[\sqrt{x^2}=?\]

OpenStudy (anonymous):

(xx)

OpenStudy (asnaseer):

no - let me try with actual numbers. what does this equal?\[\sqrt{4}=?\]

OpenStudy (anonymous):

2

OpenStudy (asnaseer):

correct, and what about this?\[2^2=?\]

OpenStudy (anonymous):

4

OpenStudy (asnaseer):

good, so therefore what will this equal?\[\sqrt{2^2}=?\]

OpenStudy (anonymous):

would it be 2 since 2X2 is 4 but the SR would be 2

OpenStudy (asnaseer):

correct, so now use that knowledge to work this out:\[\sqrt{x^2}=?\]

OpenStudy (anonymous):

x

OpenStudy (asnaseer):

correct. now similarly:\[\sqrt[3]{x^3}=x\]because the "cube root" of x "cubed" just gets you back at x. agreed?

OpenStudy (anonymous):

yes agree so then the problem would look like this \[\sqrt[3]{27+x}\] right

OpenStudy (asnaseer):

no - we are not ready to tackle that yet.

OpenStudy (anonymous):

oh k no wonder I get lost I want to finish to fast

OpenStudy (asnaseer):

in your equation you have:\[\sqrt[3]{27 (x+3)^3}\]can you think of some number that, when it is cubed, will give you 27?

OpenStudy (anonymous):

would that be like 9*3

OpenStudy (asnaseer):

another way of phrasing that would be:\[\sqrt[3]{27}=?\]it has to equal ONE number

OpenStudy (anonymous):

I believe that is the only number that goes into 27

OpenStudy (anonymous):

so then 9

OpenStudy (asnaseer):

\[9\times9\times9\ne27\]

OpenStudy (anonymous):

no that would be 3 3*3*3

OpenStudy (asnaseer):

correct, so we can rewrite your expression as:\[\sqrt[3]{27 (x+3)^3}=\sqrt[3]{3^3 (x+3)^3}\]correct?

OpenStudy (anonymous):

yes correct

OpenStudy (asnaseer):

now next thing to notice is that:\[a^3b^3=(ab)^3\]therefore:\[3^3(x+3)^3=(3(x+3))^3\]

OpenStudy (anonymous):

so then would mulitply the 3 timesx and the 9?

OpenStudy (anonymous):

no no no the 3*x and 3*3

OpenStudy (asnaseer):

no - first do you agree with that last step?

OpenStudy (anonymous):

would it be 3(x+x+3) because you have 3^3

OpenStudy (asnaseer):

I will assume you agree and proceed to the next step. we now have:\[\sqrt[3]{27 (x+3)^3}=\sqrt[3]{3^3 (x+3)^3}=\sqrt[3]{(3(x+3))^3}\]agreed?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you make it look so easy

OpenStudy (asnaseer):

:) now let us substitute:\[y=3(x+3)\]then we end up with:\[\sqrt[3]{27 (x+3)^3}=\sqrt[3]{3^3 (x+3)^3}=\sqrt[3]{(3(x+3))^3}\]\[=\sqrt[3]{y^3}\]agreed?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

and by now you should know what:\[\sqrt[3]{y^3}=?\]

OpenStudy (anonymous):

y

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

so final answer is:\[\sqrt[3]{27 (x+3)^3}=\sqrt[3]{3^3 (x+3)^3}=\sqrt[3]{(3(x+3))^3}=3(x+3)\]

OpenStudy (asnaseer):

since we said:\[y=3(x+3)\]

OpenStudy (anonymous):

wow see you did make it look easy and understandable thanks so much, can you help me with the other 2 that I have

OpenStudy (asnaseer):

<--- just post each question separately in the list on the left that way, others can also help you out - there are lots of very good helpers on this site. I may be busy with some moderator duties but I will keep an eye out for your problems.

OpenStudy (anonymous):

k thanks again

OpenStudy (asnaseer):

yw :)